OpenStudy (anonymous):

A tire dealer sells 2 tires for as much as he paid for 3 tires. He paid $240 for a dozen tires. What is the total profit on 12 tires?   $120 $180 $240 $360 $385

OpenStudy (anonymous):

240/12 = 20 <---- price paid for each tire sells 2 tires (costed him 40) for the price of 3 tires ($60) 60-40 = 20 <--- per two tires he sells 20/2 <----- 10 dollars profit per tire 12*10 = $120 <------ total profit